# HE-LRM: Encrypted Deep Learning Recommendation Models using Fully Homomorphic Encryption

### v1

@@ 2.1 CKKS
CKKS [26, 27] is a SIMD style homomorphic encryption scheme that encrypts a vector of complex or real values into a ciphertext. This scheme relies on Ring Learning with Errors (RLWE) ciphertexts in ${\mathcal{R}_{Q}^{2}}$ for a given ring $\mathcal{R}=\mathbb{Z}[X]/(X^{N}+1)$, where $N=2^{k}$, for $k>0$, is the ring dimension and $\mathcal{R}_{Q}=\mathcal{R}/Q\mathcal{R}$ describes the ring $\mathcal{R}$ reduced modulo an integer $Q=\prod_{i=0}^{L}q_{i}$. The integer $L$ is known as the multiplicative depth and represents the maximum number of rescaling levels available before decryption fails. CKKS supports element-wise operations such as addition and multiplication as well as cyclic rotation. Based on the characteristics of the CKKS scheme, it becomes a natural choice for applications such as deep learning recommendation models, which accept real-valued vectors as input. We now describe some of the core operations used with CKKS.
@@
Encoding. Consider a real (or complex) vector $\textbf{x}\in\mathbb{C}^{N/2}$. This vector can be encoded into elements of $\mathcal{R}_{Q}$ using an approximate inverse of a scaled complex canonical embedding. More precisely, one applies an inverse Fast Fourier Transform on the elements of x and scales each output by a scaling factor $\Delta$. Finally, each element is rounded to the nearest integer as encryption is performed over integers modulo $Q$. This step outputs a plaintext polynomial $\textbf{m}(X)$ which packs $N/2$ complex values. These $N/2$ values are referred to as the available slots in a CKKS polynomial.
Figure 2: Latencies (single-threaded) of primitive homomorphic operations as a function of level averaged over 30 runs. Both ciphertext-ciphertext multiplication and ciphertext rotation require a compute and memory intensive key-switching operation. Multiplication time includes rescaling (for both cases) and key-switching (for CT-CT multiplication). Due to rescaling, CKKS multiplications require at least two limbs for both operands.
@@
Rotation. Cyclic rotations shift the elements of the input vector x by $20<k<N/$ 2 slots. The resulting ciphertext after this operation corresponds to the same underlying vector with shifted elements by $k$ slots.
@@
Bootstrapping. In CKKS, a ciphertext can be bootstrapped to increase its number of available levels. This operation is necessary after the moduli chain has been depleted ($\ell=0)$ via multiplications and no more levels are available for further computations. Among all CKKS operations, bootstrapping is the most computationally expensive, taking roughly 20 seconds on a single-threaded CPU [28].
Figure 2 illustrates the latencies of the main CKKS homomorphic operations for both plaintexts (PT) and ciphertexts (CT) in terms of the level $\ell$.
2.2 Deep Learning Recommendation Model
Deep Learning Recommendation Models (DLRM) [2] are an example of a neural network-based recommendation model introduced by Facebook Research to handle both categorical (sparse) and continuous (dense) features. In order to process both types of inputs, the DLRM architecture combines multi-layer perceptrons (MLP) to process the continuous features and transforms categorical features into continuous embeddings through learned lookup tables. All processed inputs are then combined with some interaction operation. The model outputs a probability which corresponds to the likelihood of a user clicking an ad, a quantity referred to as click-through rate. More precisely, the DLRM architecture performs the following steps:
  1. 1.
  2. 2.
    The sparse (categorical) features are transformed into dense embedding vectors. One embedding table per categorical feature is used to retrieve these vectors.
  3. 3.
  4. 4.
    The concatenated features are then then fed into a top MLP that consists of a series of linear layers with ReLU activations.
  5. 5.
    The output of the top MLP is passed through a sigmoid function to produce a click-through probability.
We refer to [2] for more details about the algorithm and https://github.com/facebookresearch/dlrm. for the open-source implementation in both PyTorch and Caffe2. While originally developed for click-through rate prediction in targeted advertising, the DLRM architecture is flexible and can be adapted to process any input which contains both categorical and dense features and produce probability estimates for a wide range of applications. In this paper, we additionally apply the DLRM architecture to the the UCI Heart Disease dataset.
3 Threat Model
We assume a semi-honest threat model  [29] in which an adversary faithfully takes part in the private inference protocol but may try to learn additional information from the messages they receive. This threat model (also known as honest-but-curious) is also assumed in a majority of prior works in private inference [30, 31, 32, 33, 34, 18].
We also assume that the client knows the input dimension to the neural network for which they query. In the context of CKKS and especially private CNN inference, this assumption has been implicit: the client encrypts their image into the slots of a CKKS ciphertext before sending them to server to be processed. The ordering of the pixels within the encrypted slots must be known by the server in order for the server to perform the encrypted computation correctly.
In this paper, we make this assumption explicit given that the inputs are no longer only dense features but also the sparse features used for the embedding tables. To this end, we assume that the client knows the input size to the Bottom MLP and the sizes (number of rows) of each embedding table. Crucially, the client does not know the embedding dimension (number of columns) or parameters of the embedding tables, the interaction method, or the parameters of the MLPs.
4 The Orion Framework
In this section, we explain our rationale for choosing the Orion framework and discuss our implementation of DLRM within it. Orion is a tool that translates deep neural networks written in PyTorch [35], a widely used deep learning library, into efficient FHE programs. Its high-level interface abstracts and automates the programming complexities that traditionally complicate FHE neural inference. This automation allows researchers to quickly iterate on ideas without dealing with low-level FHE implementation details. For most feed-forward neural networks, Orion automates SIMD data packing for linear layers, bootstrap placement, and scale management. Orion also handles polynomial approximations of activation functions and dynamically manages evaluation keys by loading and storing them as needed. We discuss data packing and bootstrap placement below.

### v2

@@ Appendix A CKKS operations
CKKS cheon2017homomorphic; ckks2 is a SIMD style homomorphic encryption scheme that encrypts a vector of complex or real values into a ciphertext. This scheme relies on RingLWE ciphertexts in $\mathcal{R}_{Q}^{2}$} for a given ring $\mathcal{R}=\mathbb{Z}[X]/(X^{N}+1)$, where $N=2^{k}$, for $k>0$, is the ring dimension and $\mathcal{R}_{Q}=\mathcal{R}/Q\mathcal{R}$ describes the ring $\mathcal{R}$ reduced modulo an integer $Q=\prod_{i=0}^{L}q_{i}$. The integer $L$ is known as the multiplicative depth and represents the maximum number of rescaling levels available before decryption fails. CKKS supports element-wise operations such as addition and multiplication as well as cyclic rotation. Based on the characteristics of the CKKS scheme, it becomes a natural choice for applications such as deep learning recommendation models where real-valued vectors are considered. We now describe some of the core operations used with CKKS.
@@
Encoding. Consider a real (or complex) vector $\textbf{x}\in\mathbb{C}^{N/2}$. This vector can be encoded into elements of $\mathcal{R}_{Q}$ using an approximate inverse of a scaled complex canonical embedding. More precisely, one applies an inverse Fast Fourier Transform on the elements of x and scales each output by a scaling factor $\Delta$. Finally, each element is rounded to the nearest integer as encryption is performed over integers modulo $Q$. This step outputs a plaintext polynomial $\textbf{m}(X)$ which packs $N/2$ complex values. These $N/2$ values are referred to as the available slots in a CKKS object.
@@
Rotation. Cyclic rotations shift the elements of the input vector x by $0<k<N/2$ slots. The resulting ciphertext after this operation corresponds to the same underlying vector with shifted elements by $k$ slots.
@@
Bootstrapping. In CKKS, a ciphertext can be bootstrapped to increase its number of available levels. This operation is necessary after the moduli chain has been depleted ($\ell=0)$ via multiplications and no more levels are available for further computations. Among all CKKS operations, bootstrapping is the most computationally expensive, taking roughly 20 seconds on a single-threaded CPU.
Appendix B The HE-LRM architecture
Figure 10 illustrates the entire pipeline of our HE-LRM architecture.
Appendix C Compression ratios
Table 1 reports the compression ratios, slots required and resource requirements for different embedding table thresholds.
Table 1: Compression ratio and resource requirements for different embedding table thresholds.